# MoColl: Agent-Based Specific and General Model Collaboration for Image Captioning

### v1

@@ B.3 Our Methods
B.3.1 Hyperparameters
As for the model collaboration framework, we set the number of the few-shot examples as 5, and the maximum number of questions an agent is allowed to ask as 6 in the intra-dataset evaluation setting and 3 in the inter-dataset generalization setting. In all experiments, except those specifically stated, we always use the RAG as the ICL strategy for the agent. Notice that we can still use few-shot examples with the MIMIC-CXR training dataset in the setup of inter-dataset generation. For the generating setting, we set the max length of tokens as 4096 and temperature as 0 for both VQA model and LLM-based agent. Other hyperparameters remain the default values.
In the warm-up stage, we keep both the visual encoder and LM weights frozen, and only update the projection network. We only train 1 epoch with auto-regressive loss. We set the learning rate as 2e-3 and use cosine learning rate scheduler with warmup ratio as 0.03. The max length of tokens is 4096, and the batch size is 16 per device. Other hyperparameters remain the default values.
In the agent-guided tuning stage, we show the implementation details of the inference step, selection step, and training step individually as follows. For the inference step, we also set the max length of tokens as 4096 and temperature as 0.1 for both VQA model and LLM-based agent. The few-shot examples are captions randomly selected from the training set. For the selection step, we set the temperature of the selection agent as 0. For the training step, the visual encoder weights are still frozen, and the LM and the projection network can be updated. We train 5 epochs for the fine-tuning setting and only 1 epoch for the zero-shot setting to prevent overfitting. The loss function is auto-regressive loss with weighted KL penalty to enhance stability. We set the learning rate as 3e-7 and use cosine learning rate scheduler with warmup ratio as 0.03. The max length of tokens is 4096, the batch size is 4 per device and the gradient accumulation steps are 2. Other hyperparameters remain the default values.
@@ B.4.1 Specialized models
For all specialized models, we leverage their official code repository to reproduce the results based on our setup. Especially, for XProNet (Wang et al., 2022a) and M2KT (Yang et al., 2023), we follow their preprocessing process and use CheXbert (Smit et al., 2020) to rebuild labels with reports under our setup. We set the maximum sequence length as 530, which is the maximum length of captions in our experiments. Other hyperparameters remain default values.
@@ B.4.2 General VLMs
We use official implementations and their released model checkpoints for LLaVA-Med and LLaVA-v1.5-7B, and official APIs for Qwen2-VL-7bB-Instruction and Deepseek-VL2. The aligned VLM is trained through the warm-up stage. The fine-tuned VLM is trained on the image-instruction-caption tuple with auto-regressive loss. The visual encoder weights are still frozen, and the LM and the projection network can be updated.
For the generating setting, we provide 5 few-shot examples of captions randomly to enhance the understanding of the specific domain. We also set the temperature as 0 and other hyperparameters remain the default values.

### v3

@@ B.3 Our Methods
MoColl framework.
For the MoColl framework, we set the maximum number of questions an agent is allowed to ask as 6 in the within-dataset evaluation and 3 in the cross-dataset generalization. In all experiments, except those specifically stated, we always use the RAG as the ICL strategy for the agent. We always provide 5 few-shot examples of captions for the agent. Notice that we can still use few-shot examples with the MIMIC-CXR training dataset in the setup of cross-dataset generation. we always provide 5 few-shot examples of captions to enhance the understanding of the specific domain. For the generation setting, we set the max length of tokens as 4096 and temperature as 0 for both VQA model and LLM-based agent.
Warm-up stage.
In the warm-up stage, we keep both the visual encoder and LM weights frozen, and only update the projection network. We only train 1 epoch with auto-regressive loss. We set the learning rate as 2e-3 and use cosine learning rate scheduler with warm-up ratio as 0.03. The max length of tokens is 4096, and the batch size is 16 per device.
Agent-guided tuning stage.
In the agent-guided tuning stage, we show the implementation details of the inference step, selection step, and training step individually as follows. For the inference step, we also set the max length of tokens as 4096 and temperature as 0.1 for both VQA model and LLM-based agent. The few-shot examples are captions randomly selected from the training set. For the selection step, we set the temperature of the selection agent as 0. We can also improve the questioning prompt according to the conversation manually. For the training step, the visual encoder weights are still frozen, and the LM and the projection network can be updated. We train 5 epochs in the within-dataset evaluation and only 1 epoch in the cross-dataset generalization to prevent overfitting. The loss function is auto-regressive loss with weighted KL penalty to enhance stability. We set the learning rate as 3e-7 and use cosine learning rate scheduler with warm-up ratio as 0.03. The max length of tokens is 4096, the batch size is 4 per device and the gradient accumulation steps are 2.
Other hyperparameters remain their default values.
@@ B.4.1 Specialized models
For all specialized models, we leverage their official code repository to reproduce the results based on our setup. Especially, for XProNet Wang et al. (2022a) and M2KT Yang et al. (2023), we follow their preprocessing process and use CheXbert Smit et al. (2020) to rebuild labels with reports under our setup. We set the maximum sequence length as 530, which is the maximum length of captions in our experiments. Other hyperparameters remain default values.
@@ B.4.2 General VLMs
We use official implementations and released model checkpoints for LLaVA-Med and LLaVA-v1.5-7B, and official APIs for Qwen2-VL-7B-Instruction and Deepseek-VL2. The aligned VLM is trained through the warm-up stage. The fine-tuned VLM is trained on the image-instruction-caption tuple with auto-regressive loss. Its visual encoder weights are still frozen, and the LM and the projection network can be updated. We set its learning rate as 3e-7 and use cosine learning rate scheduler with warm-up ratio as 0.03.
For the generation setting, we provide 5 few-shot examples of captions randomly to enhance the understanding of the specific domain. We also set the temperature as 0 and other hyperparameters remain the default values.
B.4.3 MoColl framework with VQA models
We also use official implementations and released model checkpoints for LLaVA-Med and LLaVA-v1.5-7B, serving as VQA tool models instead of general VLMs working independently.